OpenStudy (anonymous):

Find the equation of the curve which has gradient function 5x^4 – 6x^2 + 4 and passes through the point (1, 8).

OpenStudy (anonymous):

I got x^5 - 2x^3 + 4x, BUT my main question in working out C, do the indicies matter? As i got 5

OpenStudy (anonymous):

8 = 1 - 2 + 4 to 8 = 3 + C C = 8 - 3 = 5 , but why dont the indicies matter? Do I always ignore them?

OpenStudy (amistre64):

y = x^5 - 2x^3 + 4x + C, since they tell you what x and y equal with the point ... (1,8)

OpenStudy (amistre64):

what indices are you talking about?

OpenStudy (amistre64):

y = x^5 - 2x^3 + 4x + C 8 = (1)^5 - 2(1)^3 + 4(1) + C 8 = 1 - 2 + 4 + C 8 = 3 + C 5 = C therefore y = x^5 - 2x^3 + 4x + 5